OpenStudy (anonymous):

how Could I find the two missing angle measures if I know some of the side lengths of a right triangle?

OpenStudy (superdavesuper):

if u got 2 sides of a right-angled triangle, then u can find the third one by c^2=a^2+b^2 then with all sides known, u can solve for the angles.

OpenStudy (anonymous):

im confused. How would that work????

OpenStudy (superdavesuper):

which sides are given in the right-angled triangle?

OpenStudy (anonymous):

Im looking for angles and need to solve with lengths. and I have no idea. there isnt a certain problem

OpenStudy (superdavesuper):

drawing tool is down in OS :( but u know in a right-angled triangle, c^2=a^2+b^2 where c is the longest side right?

OpenStudy (anonymous):

yeah. the hypotenuse?

OpenStudy (superdavesuper):

yup yup....so if u get any 2 side lengths, u can solve for the missing one ok?

OpenStudy (anonymous):

yeah but I thought if i were looking for an angle i could solve with sides?

OpenStudy (superdavesuper):

yup...u can then either the trig ratios such as sinA=a/c or cosA=b/c to find the angles.

OpenStudy (superdavesuper):

u can then USE either...

OpenStudy (anonymous):

what do I do after I have what I have : cos(x degrees) = adjacent/hypotenuse cos(x degrees) = 23/25 cos(x degrees) = 0.92

OpenStudy (superdavesuper):

then u can use arcos or cos^-1 or inv.cos on ur calculator to find the angle

OpenStudy (anonymous):

this doesnt look right.. ≈0.402715842

OpenStudy (superdavesuper):

hehehe thats because ur calculator is set to "radian" - look for a button that said "degree" it should come out to about 23 deg.

OpenStudy (anonymous):

ok I was just using a bad online calculator. Thank you so much!!!!

OpenStudy (superdavesuper):

welcome :)